OpenStudy (anonymous):

Help I need to solve for y using addition method x+4y=37 8x+4y=100

OpenStudy (anonymous):

x+4y=37 8x+4y=100 ____________- -7x + 0 = -63 x = 9 now fill in x to get y

OpenStudy (anonymous):

x+4y=37 ----> -8x - 32y = -296 8x+4y=100 -----> 8x + 4y = 100 ------------- + -28 y = - 196 y = 7

OpenStudy (anonymous):

thank you